OpenStudy (anonymous):

lim for x to - infinity of x-log(2^(-x)+1)

OpenStudy (anonymous):

\[\lim_{x\rightarrow -\infty}x-\log(2^{-x}+1)\]

OpenStudy (anonymous):

it's that how can you handle with it?

OpenStudy (anonymous):

oh sorry there's a plus instead of a min in front of the log

OpenStudy (anonymous):

i was just writing to see if this was the right thing i guess it is maybe \[\lim_{x\rightarrow -\infty}x+\log(2^{-x}+1)\]??

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

sorry browser took a dump

OpenStudy (anonymous):

there might be some fancy way to do this that i do not know, but i would think of it like this. the +1 is clearly irrelevant and if you had \[x-\log(2^{-x})\] this is the same as \[x+x\log(2)\] and it is pretty clear that this goes to - infinity as x does

OpenStudy (anonymous):

ok that lousy answer bothered me so lets try this \[\log(2^{-x}+1)=\log(\frac{1}{2^x}+1)\] \[=\log(\frac{1+2^x}{2^x})=\log(1+2^x)-\log(2^x)=\log(1+2^x)-x\log(2)\] and so \[x+\log(2^{-x}+1)=x-x\log(2)+\log(1+2^x)\] now take the limit as x goes to minus infinity

OpenStudy (anonymous):

since \[\log(2)<1\] we know \[1-\log(2)> 0\] and so \[\lim_{x\rightarrow -\infty} (1-\log(2))x=-\infty\] and also \[\lim_{x\rightarrow -\infty} \log(1+2^x)=0\] and therefore we get \[-\infty\]

OpenStudy (anonymous):

thank you very much you have been very helpfull